A man did a job for $360. It took him 6 hours longer than he expected and therefore he earned $2 per hour less than he anticipated. How long did he expect that it would take to do the job?
Answer by stanbon(48568)

A man did a job for $360. It took him 6 hours longer than he expected and therefore he earned $2 per hour less than he anticipated. How long did he expect that it would take to do the job?
--------
Let the time he expected to use be "x" hrs.
-------------------------------------------------
He received 360/(x+6) dollars per hour.
---------------
He expected to make 360/x dollars per hour
-------------------
Equation:
Expected rate - received rate = 2 dollars
360/x - 360/(x+6) = 2
Multiply thru by x(x+6) to get:
360(x+6) - 360x = 2x(x+6)
2x^2 + 12x - 360*6 = 0
x^2 + 6x - 1080 = 0
(x-30)(x+36) = 0
Positive solution:
x = 30 hrs (amount of time he expected to work to do the job)
Cheers,
Stan H.